A box contains 8 green, 9 red, 5 yellow, and 7 blue marbles. Four marbles are drawn from the box at the same time.
a) What is the probability that all four marbles are the same color?
b) What is the probability that none of the marbles is red?
c) What is the probability that you get one marble of each color?
d) What is the probability that you draw more than 2 yellow marbles?

Answer by ewatrrr(24785)

 
Hi,
29 marbles: 8 green, 9 red, 5 yellow and 7 blue
4 drawn at the same time
P(all four marbles are the same color) = 8C4/29C4 + 9C4/29C4 + 5C4/29C4 + 7C4/29C4
P(none of the marbles is red) = 1 - P(all red) = 1 - 9C4/29C4
P(one marble of each color) = 8C1*9C1*5C1*7C1/29C4
P(more than 2 yellow marbles) = 29C1*5C3/29C4 + 5C4/29C4